OpenStudy (anonymous):

Convert (-5, pi/4) to rectangular coordinates. A. (sqrt(2) / 2, sqrt(2) / 2) B. (5 sqrt(3) / 2, 5/ 2) C. (5 sqrt(2) / 2, 5 sqrt(2) / 2) D. (-5 sqrt(2) / 2, -5 sqrt(2) / 2)

OpenStudy (anonymous):

I am thinking that it would be B. However I am having trouble understanding how to get my answer.

OpenStudy (anonymous):

\[x = rcos \theta\]\[y= rsin \theta\]

OpenStudy (anonymous):

\[x = (-5)\cos(\frac{ \pi }{ 4 })\]\[y = (-5)\sin(\frac{ \pi }{ 4 })\]

OpenStudy (anonymous):

I don't get any of the answers :(

OpenStudy (anonymous):

yes. last step is to use a unit circle to get sin (π/4) and cos (π/4)

OpenStudy (anonymous):

45?

OpenStudy (anonymous):

Oh sin and cos of 45?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It would be A?

OpenStudy (anonymous):

D. A doesn't have -5 for r

OpenStudy (anonymous):

Oh you just said the answer didn't you

OpenStudy (anonymous):

yes r can be negative.